The ship was moved 0.9 nmi to the GPS coordinates for Site 1005, and Hole 1005A was initiated at 2000 hr, 7 March. The water depth was determined to be 350.7 mbsl based on recovery of the mudline core. APC Cores 1005A-1H to 9H were cut from 0 to 67.5 mbsf with 101.3% recovery, orienting Cores 3H-9H with the tensor tool. Adara temperature tool heat flow measurements were deferred until Hole 1005B to determine the approximate depth of hard layers that might damage the tool. As in the previous two sites, recovery was reduced by hard layers jamming in the throat of the bit. Core liners shattered on Cores 8H and 9H, prompting a change to the XCB coring system. XCB Cores 10X to 53X were cut from 67.5 to 462.4 mbsf with 20.6% recovery. Cores 14X to 29X and 37X to 53X had only 12% recovery in soft, gassy, sediments with lithoclasts and occasional cemented layers, despite using minimal circulation rates. The pipe became stuck while cutting Core 38X at 686.2 mbsf, when the hole apparently collapsed, packing off the annulus and stalling the rotary. The formation was fractured and the drill string was eventually worked free after circulation and rotation were re-established. Subsequently, the hole was back-reamed to 172 mbsf, and then reamed to bottom. WSTP temperature probes were deployed after Cores 11X, 14X, 19X, and 22X. As a result of the unstable hole conditions, a precautionary short trip was made to 255.6 mbsf to condition the hole prior to logging.
Because of the previous difficulty logging Hole 1003C, we chose to log Hole 1005A rather than risk not obtaining any logging data, if similar unstable hole conditions were encountered in Hole 1005C. Therefore, XCB coring was terminated at 462.4 mbsf as the result of poor recovery. The bit was pulled up to the logging depth at 103.8 mbsf. A Sonic/DIT log was run to 445 mbsf, and an IPLT log was run to 434 mbsf. Finally, a VSP log was run to 409 mbsf. After running the VSP experiment in the open hole, the VSP tool would not reenter the bit because the hydraulic clamping arm would not retract; therefore, a Kinley crimper and cutter tools were dropped to clamp the line inside the BHA and then sever the logging line above the tool. The VSP tool was retrieved intact and the pipe cleared the rotary at 2030 hr on 10 March.
Coring at Hole 1005B was initiated 10 m to the south of Hole 1005A. The hole was spudded at 0011 hr on 11 March using the same APC BHA without the nonmagnetic drill collars (NMDC). Hole 1005B was drilled primarily to obtain heat-flow measurements. The water depth was determined to be 351.7 mbsl based on recovery of the mudline core. APC Cores 1005B-1H to 7H were cut from 0 to 51.0 mbsf with 81.4% recovery. Adara temperature tool heat-flow measurements were taken during coring of Cores 3H through 7H. Core 7H was a partial stroke that threw slack in the coring wireline where it entangled on the heave compensator and the line subsequently parted. After restringing the wireline, Core 7H was retrieved. XCB Core 8X was taken from 51.0 to 52.0 mbsf to penetrate a hard layer, and XCB Core 9X was cut from 52.0 to 61.7 mbsf. A WSTP temperature probe was deployed after Cores 8X and 9X. With the objectives of Hole 1005B achieved, the hole was terminated and the bit cleared the rotary table at 1305 hr on 11 March.
The ship was moved 20 m south of Hole 1005B, and Hole 1005C was initiated with a RCB BHA with a mechanical bit release (MBR) at 1750 hr on 11 March. The water depth was estimated to be 351.6 mbsl based on bit contact. A hole was drilled with a center bit to 386.6 mbsf. RCB Cores 1R to 34R were cut from 386.6 to 700.0 mbsf with 34.1% recovery.
A conditioning trip for logs was made to 85.5 mbsf and the bit was released using the MBR, the hole was filled with sepiolite mud, and the open drill string was positioned at 398.3 mbsf. The following suite of logs were run: (1) an IPLT log to 700.0 mbsf; (2) a DIT/Sonic log to 693.0 mbsf; (3) a FMS to 613 mbsf; and (4) a VSP log to 541 mbsf. The FMS and VSP logs were lighter weight tools that limited the ability to work them deeper into the hole. After plugging the hole with cement and gel mud, the pipe was pulled and secured for sea voyage at 0854 hr on 15 March.
Site 1006 Operations Report
166 Table of Contents